# Higher Order Dualities between Prime Ideals

### v1

@@ 2 General Duality Between Prime Ideals
Thus, along with a negative sign, the sum counts the number of prime ideals containing the ideal $I$ with the second largest norm. Hence, combining (2.2), (2.6) and (2.10), we conclude our theorem. ∎
Therefore, we have what we call the second order duality between prime ideals in a number ring. We note that the RHS of Lemma 2.2 counts the number of prime ideals with the largest norm and that of Theorem 2.3 counts the number of prime ideals with the second largest norm. Hence, we call them first and second order dualities respectively. Let us now state and prove the general higher order dualities between prime ideals. The proof will follow an idea similar to that used in the proof of Theorem 2.3
@@
$\sum_{J\supset I}\mu_{K}(J){{\omega_{K}(J)-1}\choose k-1}f(J)=(-1)^{k}Q_{C}^{k}(I)$
@@
Apriori, we must note that this theorem only makes sense when the chosen ideal $I$ has prime factors with at least $k$ distinct norms. We now start with our usual factorization of $I$ as done above: $I=I_{1}...I_{s}$ where the $I_{i}$’s are product of prime ideals, containing $I$, of the same norm. Without the loss of generality, let us assume that the factors are written in an ascending order of their norms, i.e. $N(I_{i})<N(I_{j})$, for all $i<j$. From the LHS of the above equation, we have
@@
$=$ $\frac{1}{(k-1)!}\sum_{j=1}^{s-k+1}\sum_{\begin{subarray}{c}J\supset I_{j}\\ J\not\supset I_{k^{\prime}}\forall k^{\prime}\geq j+1\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})(\omega_{K}(J)+\omega_{K}(J^{\prime})-1)...(\omega_{K}(J)+\omega_{K}(J^{\prime})-k+1)$ (2.11)
We note that the summand of the innermost sum in (2.11) can be repersented as
$\sum_{u=0}^{k-1}\mu_{K}(J^{\prime})c_{u}(\omega_{K}(J))\omega_{K}(J^{\prime})^{u}$ (2.12)
@@
$\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})(\omega_{K}(J)+\omega_{K}(J^{\prime})-1)...(\omega_{K}(J)+\omega_{K}(J^{\prime})-k+1)=\sum_{u=0}^{k-1}c_{u}(\omega_{K}(J))\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})^{u}$ (2.13)
@@
$\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})^{u}=\sum_{i=1}^{t}(-1)^{i}{t\choose i}i^{u}$ (2.14)
@@
$\sum_{\ell=1}^{u}\tilde{c}_{\ell}(t)x^{\ell}(1-x)^{t-\ell}=\sum_{i=1}^{t}(-1)^{i}{t\choose i}i^{u}x^{i}$ (2.18)
With $x=1$, the RHS of (2.18) is exactly the RHS of (2.14), while the LHS of (2.18) with $x=1$ is non-zero only when $\ell=t$. That is only possible if $t\leq u$. Now, $t$ is the number of distinct prime ideals containing the product $I_{j+1}...I_{s}$, where $j\leq s-k+1$. Hence, by our hypothesis, $k-1\leq t$. But $u\leq k-1$, and therefore, the only non-zero contribution occurs by RHS of (2.13) is when u=k1u=k-1. In other words, we can also say that for $0\leq u\leq k-2$, we have that
$\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})^{u}=0$ (2.19)
Thus, from (2.11), (2.13) and (2.19), we continue to get
@@
$=$ $\frac{1}{(k-1)!}\sum_{j=1}^{s-k+1}\sum_{\begin{subarray}{c}J\supset I_{j}\\ J\not\supset I_{k^{\prime}}\forall k^{\prime}\geq j+1\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})^{k-1}$ (2.20)
@@
$=$ $\frac{1}{(k-1)!}\sum_{j=1}^{s-k+1}\sum_{\begin{subarray}{c}J\supset I_{j}\\ J\not\supset I_{k^{\prime}}\forall k^{\prime}\geq j+1\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})(\omega_{K}(J^{\prime})-1)...(\omega_{K}(J^{\prime})-k+2)$ (2.21)
Again the inner sum of (2.21) can be written as
@@
$=$ $\begin{cases}(-1)^{k-1}(k-1)!,\;\text{when $I_{j+1}...I_{s}$ has exactly $k-1$ distinct prime factors}\\ 0,\;\text{otherwise}\end{cases}$ (2.22)
@@
$\sum_{J\supset I}\mu_{K}(J){{\omega_{K}(J)-1}\choose k-1}f(J)=\frac{(-1)^{k-1}(k-1)!}{(k-1)!}\sum_{\begin{subarray}{c}J\supset I_{s-k+1}\\ J\not\supset I_{k}^{\prime}\forall k^{\prime}\geq s-k\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)=(-1)^{k}Q_{C}^{k}(I)$ (2.23)
@@
In the following section, we will prove estimates of functions counting prime ideals that will be essential for us in proving our main results that will follow later, where also, the above duality identities will be used predominantly.
@@ 3 The Prime Ideal Counting Functions and their Estimates
$\rho(\beta)$ $=1,\;\text{for $0\leq\beta\leq 1$}$
$-\beta\rho^{\prime}(\beta)$ $=\rho(\beta-1),\;\text{for $\beta>1$}$
@@
Tenenbaum [Te00] has also proved several estimates involving the $\Psi(x,y)$ function and the $k^{th}$ largest prime factors which proved to be very useful in proving results in the classical case. But here, as we shift from the classical to an algebraic setting, we need to consider the algebraic analogues of such functions and their corresponding estimates. We count, what are known as smooth ideals, using the following function: for a number field $K$,
$\Psi_{K}(X,Y)=\sum_{\begin{subarray}{c}N(I)\leq X\\ M_{1}(I)\leq Y\end{subarray}}1$ (3.2)
counts the number of ideals in $\mathcal{O}_{K}$ with their norm bounded by $X$ such that the norm of their prime factors are bounded by $Y$. It has been proved by Krause [Kr90] and Moree [Mo92] that an asymptotic estimate for the function $\Psi_{K}$ exist and is given by
@@
where $\beta=\frac{\log X}{\log Y}$. Note that this asymptote is uniform for $1\leq\beta\leq(\log X)^{1-\varepsilon}$. (3.3) also corresponds to Hildebrand’s result [Hi86] in the classical case, i.e. when $K=\mathbb{Q}$. The implicit constant in the error of (3.3) depends on $\varepsilon$ which can be suitably chosen to derive desired estimates. As done in the classical case by Maier [Ma(up)] as a corollary to Hildebrand’s result, in the algebraic setting too, one can use the following upper bound of the Dickman function (see Norton [No71])
@@
$\rho(\beta)\sim\frac{1}{\sqrt{2\pi\beta}}e^{-\beta\log\left(\frac{\beta}{e}\right)}$ (3.4)
and then use (3.4) to get another useful estimate of $\Psi_{K}(X,Y)$, i.e.
@@
for some constant $c_{3}$. We will use both the estimates (3.3) and (3.5) in different contexts in our proofs. As mentioned earlier, our aim is to derive new formulas of the Chebotarev Density using higher order dualities between prime ideals as proved above. In this paper, we will mainly focus on using the second order duality777can be viewed as generalizations of results by Alladi-Johnson [AJ24] and the author [Se25] and comment on how the higher order dualities can be utilized following in lines of the proofs by Alladi-Sengupta [AS(ip)]. In what follows, we will prove a few important lemmas in this section that we will use to prove our main results.
Looking at the hypothesis of Theorem 2.3, we observe that it is required to consider only such ideals that satisfy the condition $Q^{1}(I)=1$. We now prove that the number of ideals that do not satisfy such a condition is small compared to a bound of the norm of the ideals. In fact, in the lemma that follows, we prove something even stronger.
@@
$\mathcal{N}_{1}(K;X)\ll\frac{X}{e^{c_{3}\sqrt{\log X\log\log X}}}$ (3.7)
@@
$\Psi_{K}\left(X,e^{\sqrt{\frac{\log X}{\log\log X}}}\right)\ll\frac{X}{e^{c_{3}\sqrt{\log X\log\log X}}}$ (3.8)
@@
counts the number of ideals in $\mathcal{O}_{K}$ with norm bounded by $X$ and the second largest norm among all their prime ideal factors is bounded by $Y$. The next lemma will provide us with a crucial estimate of $\Psi_{K,2}(X,Y)$ with $Y$ belonging in some given interval. This estimate is for the number of ideals with all the prime factors, except ones with the largest norm, with of smaller norms. From the estimate (3.5), it is quite clear that $\Psi_{K}(X,Y)$ is quite small as compared to $X$, for large $X$. On the other hand, $\Psi_{K,2}(X,Y)$ is not that small. This can be viewed by considering ideals of the form $I=\mathfrak{p}_{1}\mathfrak{p}_{2}$, such that $N(\mathfrak{p}_{1})=2<N(\mathfrak{p}_{2})$. We observe that
@@
$\Psi_{K,2}(X,Y)\gg\frac{X}{\log\frac{X}{2}}$
Therefore, it is important to find its estimate, of course under some restrictions, which is provided by the following lemma.
@@
$\Psi_{K,2}(X,\mathfrak{p})=\sum_{\begin{subarray}{c}N(I)\leq X\\ M_{2}(I)=N(\mathfrak{p})\end{subarray}}1$ (3.13)
@@
$\Psi_{K,2}(X,Y)\leq\sum_{\begin{subarray}{c}\mathfrak{p}\subset\mathcal{O}_{K}\\ N(\mathfrak{p})\leq Y\end{subarray}}\Psi_{K,2}(X,\mathfrak{p})$ (3.14)
The reason behind the inequality above is that in the sum on the left, we are counting ideals, whereas on the right we are counting prime ideals with bounded norm and therefore, there might be over-counting due to the presence of an ideal inside multiple prime ideals of the same norm. Therefore, estimating the sum on the right of (3.14) would suffice for our required result.
@@
$\Psi_{K,2}(X,\mathfrak{p})=\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\sum_{N(\mathfrak{p})\leq N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{mp})}}1+O\left(\frac{X}{e^{\sqrt{\log X\log\log X}}}\right)$ (3.15)
@@
$\Psi_{K,2}(X,\mathfrak{p})\ll\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}$ (3.16)
@@
$\Psi_{K,2}(X,\mathfrak{p})\ll\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq T\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}+\sum_{\begin{subarray}{c}T<N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}$ (3.17)
@@
$\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq T\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}\ll\frac{X}{N(\mathfrak{p})\log X}\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq T\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{1}{N(\mathfrak{m})}\ll\frac{X}{N(\mathfrak{p})\log X}\prod_{\begin{subarray}{c}\mathfrak{t}\subset\mathcal{O}_{K}\\ \mathfrak{t}\text{ is prime}\\ N(\mathfrak{t})\leq N(\mathfrak{p})\end{subarray}}\left(1-\frac{1}{\mathfrak{t}}\right)^{-1}$ (3.18)
The inner product on the right can be then estimated using the Merten’s Theorem for prime ideals [Le23] as follows:
$\prod_{\begin{subarray}{c}\mathfrak{t}\subset\mathcal{O}_{K}\\ \mathfrak{t}\text{ is prime}\\ N(\mathfrak{t})\leq N(\mathfrak{p})\end{subarray}}\left(1-\frac{1}{\mathfrak{t}}\right)^{-1}\ll\log N(\mathfrak{p})$ (3.19)
@@
$\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq T\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}\ll\frac{X\log N(\mathfrak{p})}{N(\mathfrak{p})\log X}$ (3.20)
@@
$\sum_{\begin{subarray}{c}T<N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}\ll\frac{X}{N(\mathfrak{p})\log N(\mathfrak{p})}\sum_{\begin{subarray}{c}T<N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{1}{N(\mathfrak{m})}$ (3.21)
@@
$\sum_{\begin{subarray}{c}T<N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{1}{N(\mathfrak{m})}=\sum_{T<n\leq\frac{X}{N(p)^{2}}}\frac{1}{n}\sum_{\begin{subarray}{c}\mathfrak{m}\subset\mathcal{O}_{K}\\ N(\mathfrak{m})=n\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}1\leq\sum_{T<n\leq\frac{X}{N(\mathfrak{p})^{2}}}\frac{\Psi^{K}(n,N(\mathfrak{p}))}{n}\ll\sum_{T<n\leq\frac{X}{N(\mathfrak{p})^{2}}}e^{-c_{3}\alpha}$ (3.22)

### v2

@@ 2 General Duality Between Prime Ideals
Thus, along with a negative sign, the sum counts the number of prime ideals containing the ideal $I$ with the second largest norm. Hence, combining (2.2), (2.6), and (2.10), we conclude our theorem. ∎
Therefore, we have what we call the second order duality between prime ideals in a number ring. We note that the RHS of Lemma 2.2 counts the number of prime ideals with the largest norm and that of Theorem 2.3 counts the number of prime ideals with the second largest norm. Hence, we call them first and second order dualities, respectively. Let us now state and prove the general higher order dualities between prime ideals. The proof will follow an idea similar to that used in the proof of Theorem 2.3.
@@
$\sum_{J\supset I}\mu_{K}(J){{\omega_{K}(J)-1}\choose k-1}f(J)=(-1)^{k}Q_{C}^{k}(I).$
@@
A priori, we must note that this theorem only makes sense when the chosen ideal $I$ has prime factors with at least $k$ distinct norms. We now start with our usual factorization of $I$ as done above: $I=I_{1}...I_{s}$ where the $I_{i}$’s are products of prime ideals, containing $I$, of the same norm. Without the loss of generality, let us assume that the factors are written in an ascending order of their norms, i.e. $N(I_{i})<N(I_{j})$, for all $i<j$. From the LHS of the above equation, we have
@@
$=$ $\frac{1}{(k-1)!}\sum_{j=1}^{s-k+1}\sum_{\begin{subarray}{c}J\supset I_{j}\\ J\not\supset I_{k^{\prime}}\forall k^{\prime}\geq j+1\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})(\omega_{K}(J)+\omega_{K}(J^{\prime})-1)...(\omega_{K}(J)+\omega_{K}(J^{\prime})-k+1).$ (2.11)
We note that the summand of the innermost sum in (2.11) can be represented as
$\sum_{u=0}^{k-1}\mu_{K}(J^{\prime})c_{u}(\omega_{K}(J))\omega_{K}(J^{\prime})^{u}.$ (2.12)
@@
$\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})(\omega_{K}(J)+\omega_{K}(J^{\prime})-1)...(\omega_{K}(J)+\omega_{K}(J^{\prime})-k+1)=\sum_{u=0}^{k-1}c_{u}(\omega_{K}(J))\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})^{u}.$ (2.13)
@@
$\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})^{u}=\sum_{i=1}^{t}(-1)^{i}{t\choose i}i^{u}.$ (2.14)
@@
$\sum_{\ell=1}^{u}\tilde{c}_{\ell}(t)x^{\ell}(1-x)^{t-\ell}=\sum_{i=1}^{t}(-1)^{i}{t\choose i}i^{u}x^{i}.$ (2.18)
With $x=1$, the RHS of (2.18) is exactly the RHS of (2.14), while the LHS of (2.18) with $x=1$ is non-zero only when $\ell=t$. That is only possible if $t\leq u$. Now, $t$ is the number of distinct prime ideals containing the product $I_{j+1}...I_{s}$, where $j\leq s-k+1$. Hence, by our hypothesis, $k-1\leq t$. But $u\leq k-1$, and therefore, the only non-zero contribution occurs on the RHS of (2.13) when u=k1u=k-1. In other words, we can also say that for $0\leq u\leq k-2$, we have that
$\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})^{u}=0.$ (2.19)
Thus, from (2.11), (2.13), and (2.19), we continue to get
@@
$=$ $\frac{1}{(k-1)!}\sum_{j=1}^{s-k+1}\sum_{\begin{subarray}{c}J\supset I_{j}\\ J\not\supset I_{k^{\prime}}\forall k^{\prime}\geq j+1\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})^{k-1}.$ (2.20)
@@
$=$ $\frac{1}{(k-1)!}\sum_{j=1}^{s-k+1}\sum_{\begin{subarray}{c}J\supset I_{j}\\ J\not\supset I_{k^{\prime}}\forall k^{\prime}\geq j+1\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)\sum_{J^{\prime}\supset I_{j+1}...I_{s}}\mu_{K}(J^{\prime})\omega_{K}(J^{\prime})(\omega_{K}(J^{\prime})-1)...(\omega_{K}(J^{\prime})-k+2).$ (2.21)
Again, the inner sum of (2.21) can be written as
@@
$=$ $\begin{cases}(-1)^{k-1}(k-1)!,\;\text{when $I_{j+1}...I_{s}$ has exactly $k-1$ distinct prime factors}\\ 0,\;\text{otherwise.}\end{cases}$ (2.22)
@@
$\sum_{J\supset I}\mu_{K}(J){{\omega_{K}(J)-1}\choose k-1}f(J)=\frac{(-1)^{k-1}(k-1)!}{(k-1)!}\sum_{\begin{subarray}{c}J\supset I_{s-k+1}\\ J\not\supset I_{k}^{\prime}\forall k^{\prime}\geq s-k\\ J\neq\mathcal{O}_{K}\end{subarray}}\mu_{K}(J)f(J)=(-1)^{k}Q_{C}^{k}(I).$ (2.23)
@@
In the following section, we will prove estimates of functions counting prime ideals that will be essential for us in proving our main results that will follow later, where the above duality identities will be used predominantly.
@@ 3 The Prime Ideal Counting Functions and their Estimates
$\rho(\beta)$ $=1,\;\text{for $0\leq\beta\leq 1$},$
$-\beta\rho^{\prime}(\beta)$ $=\rho(\beta-1),\;\text{for $\beta>1$}.$
@@
Tenenbaum [Te00] has also proved several estimates involving the $\Psi(x,y)$ function and the $k^{th}$ largest prime factors, which proved to be very useful in proving results in the classical case. But here, as we shift from the classical to an algebraic setting, we need to consider the algebraic analogues of such functions and their corresponding estimates. We count, what are known as smooth ideals, using the following function: for a number field $K$,
$\Psi_{K}(X,Y)=\sum_{\begin{subarray}{c}N(I)\leq X\\ M_{1}(I)\leq Y\end{subarray}}1.$ (3.2)
counts the number of ideals in $\mathcal{O}_{K}$ with their norm bounded by $X$ such that the norm of their prime factors is bounded by $Y$. It has been proved by Krause [Kr90] and Moree [Mo92] that an asymptotic estimate for the function $\Psi_{K}$ exist and is given by
@@
where $\beta=\frac{\log X}{\log Y}$. Note that this asymptote is uniform for $1\leq\beta\leq(\log X)^{1-\varepsilon}$. (3.3) also corresponds to Hildebrand’s result [Hi86] in the classical case, i.e. when $K=\mathbb{Q}$. The implicit constant in the error of (3.3) depends on $\varepsilon$, which can be suitably chosen to derive desired estimates. As done in the classical case by Maier [Ma(up)] as a corollary to Hildebrand’s result, in the algebraic setting too, one can use the following upper bound of the Dickman function (see Norton [No71])
@@
$\rho(\beta)\sim\frac{1}{\sqrt{2\pi\beta}}e^{-\beta\log\left(\frac{\beta}{e}\right)},$ (3.4)
and then, use (3.4) to get another useful estimate of $\Psi_{K}(X,Y)$, i.e.
@@
for some constant $c_{3}$. We will use both the estimates (3.3) and (3.5) in different contexts in our proofs. As mentioned earlier, our aim is to derive new formulas of the Chebotarev Density using higher order dualities between prime ideals as proved above. In this paper, we will mainly focus on using the second order duality777can be viewed as generalizations of results by Alladi-Johnson [AJ24] and the author [Se25] and comment on how the higher order dualities can be utilized following in the lines of the proofs by Alladi-Sengupta [AS(ip)]. In what follows, we will prove a few important lemmas in this section that we will use to prove our main results.
Looking at the hypothesis of Theorem 2.3, we observe that it is required to consider only such ideals that satisfy the condition $Q^{1}(I)=1$. We now prove that the number of ideals that do not satisfy such a condition is small compared to a bound on the norm of the ideals. In fact, in the lemma that follows, we prove something even stronger.
@@
$\mathcal{N}_{1}(K;X)\ll\frac{X}{e^{c_{3}\sqrt{\log X\log\log X}}}.$ (3.7)
@@
$\Psi_{K}\left(X,e^{\sqrt{\frac{\log X}{\log\log X}}}\right)\ll\frac{X}{e^{c_{3}\sqrt{\log X\log\log X}}}.$ (3.8)
@@
counts the number of ideals in $\mathcal{O}_{K}$ with norm bounded by $X$, and the second largest norm among all their prime ideal factors is bounded by $Y$. The next lemma will provide us with a crucial estimate of $\Psi_{K,2}(X,Y)$ with $Y$ belonging to some given interval. This estimate is for the number of ideals with all the prime factors, except those with the largest norm, of smaller norms. From the estimate (3.5), it is quite clear that $\Psi_{K}(X,Y)$ is quite small as compared to $X$, for large $X$. On the other hand, $\Psi_{K,2}(X,Y)$ is not that small. This can be viewed by considering ideals of the form $I=\mathfrak{p}_{1}\mathfrak{p}_{2}$, such that $N(\mathfrak{p}_{1})=2<N(\mathfrak{p}_{2})$. We observe that
@@
$\Psi_{K,2}(X,Y)\gg\frac{X}{\log\frac{X}{2}}.$
Therefore, it is important to find its estimate, of course, under some restrictions, which is provided by the following lemma.
@@
$\Psi_{K,2}(X,\mathfrak{p})=\sum_{\begin{subarray}{c}N(I)\leq X\\ M_{2}(I)=N(\mathfrak{p})\end{subarray}}1.$ (3.13)
@@
$\Psi_{K,2}(X,Y)\leq\sum_{\begin{subarray}{c}\mathfrak{p}\subset\mathcal{O}_{K}\\ N(\mathfrak{p})\leq Y\end{subarray}}\Psi_{K,2}(X,\mathfrak{p}).$ (3.14)
The reason behind the inequality above is that in the sum on the left, we are counting ideals, whereas on the right, we are counting prime ideals with bounded norm, and therefore, there might be over-counting due to the presence of an ideal inside multiple prime ideals of the same norm. Therefore, estimating the sum on the right of (3.14) would suffice for our required result.
@@
$\Psi_{K,2}(X,\mathfrak{p})=\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\sum_{N(\mathfrak{p})\leq N(\mathfrak{q})\leq\frac{X}{N(\mathfrak{mp})}}1+O\left(\frac{X}{e^{c_{3}\sqrt{\log X\log\log X}}}\right).$ (3.15)
@@
$\Psi_{K,2}(X,\mathfrak{p})\ll\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}.$ (3.16)
@@
$\Psi_{K,2}(X,\mathfrak{p})\ll\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq T\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}+\sum_{\begin{subarray}{c}T<N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}.$ (3.17)
@@
$\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq T\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}\ll\frac{X}{N(\mathfrak{p})\log X}\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq T\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{1}{N(\mathfrak{m})}\ll\frac{X}{N(\mathfrak{p})\log X}\prod_{\begin{subarray}{c}\mathfrak{t}\subset\mathcal{O}_{K}\\ \mathfrak{t}\text{ is prime}\\ N(\mathfrak{t})\leq N(\mathfrak{p})\end{subarray}}\left(1-\frac{1}{\mathfrak{t}}\right)^{-1}.$ (3.18)
The inner product on the right can then be estimated using Merten’s Theorem for prime ideals [Le23] as follows:
$\prod_{\begin{subarray}{c}\mathfrak{t}\subset\mathcal{O}_{K}\\ \mathfrak{t}\text{ is prime}\\ N(\mathfrak{t})\leq N(\mathfrak{p})\end{subarray}}\left(1-\frac{1}{\mathfrak{t}}\right)^{-1}\ll\log N(\mathfrak{p}).$ (3.19)
@@
$\sum_{\begin{subarray}{c}N(\mathfrak{m})\leq T\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}\ll\frac{X\log N(\mathfrak{p})}{N(\mathfrak{p})\log X}.$ (3.20)
@@
$\sum_{\begin{subarray}{c}T<N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{X}{N(\mathfrak{mp})\log\left(\frac{X}{N(\mathfrak{mp})}\right)}\ll\frac{X}{N(\mathfrak{p})\log N(\mathfrak{p})}\sum_{\begin{subarray}{c}T<N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{1}{N(\mathfrak{m})}.$ (3.21)
@@
$\sum_{\begin{subarray}{c}T<N(\mathfrak{m})\leq\frac{X}{N(\mathfrak{p})^{2}}\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}\frac{1}{N(\mathfrak{m})}=\sum_{T<n\leq\frac{X}{N(p)^{2}}}\frac{1}{n}\sum_{\begin{subarray}{c}\mathfrak{m}\subset\mathcal{O}_{K}\\ N(\mathfrak{m})=n\\ M_{1}(\mathfrak{m})\leq N(\mathfrak{p})\end{subarray}}1\leq\sum_{T<n\leq\frac{X}{N(\mathfrak{p})^{2}}}\frac{\Psi_{K}(n,N(\mathfrak{p}))}{n}\ll\sum_{T<n\leq\frac{X}{N(\mathfrak{p})^{2}}}e^{-c_{3}\alpha},$ (3.22)